Butane, C4H10, is an easily liquefied gaseous fuel. Calculate the density of butane gas at 0.786 atm and 24°C. Give the answer in grams per liter.

Density= g/L
Answer #1

CqHro mer mass 58-1231mo P0.7.6 atm, T=242+243 = 297K. Pv-neT Reamnge Te idad ges loo MMm RT V Mm-P RT P Sedasit Mm P RT 58.1